# Flux-balance laws for spinning bodies under the gravitational self-force

### v1

@@ V.2 Spinning particles
and shows that $\Pi^{AB}$ is degenerate. However, as we remarked above in Sec. II.1, this poses no issue; in fact, we simply treat $C_{\circ}$ as an additional constraint direction with no difficulty.
@@
We now discuss the consequences of the existence of these coordinates. First, from Eqs. (144) and (146a), it follows that, instead of Eq. (2), we find a very mild generalization:
$\left\langle\dot{P}_{\alpha}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=-[A^{\beta}{}_{\alpha}(\partial_{\vartheta^{\beta}})^{A}+A_{\alpha}(\partial_{\upsilon})^{A}]\mathcal{F}_{A}.$ (147)
This is very important, as it implies that one can still write down flux-balance laws for the $P_{\alpha}$, which in this case are given by $\mu^{2}$, $E$, $L_{z}$, and the generalization of the Carter constant $K$. While such flux-balance laws have been known for $E$ and $L_{z}$ [27, 15], this shows that straightforward generalizations exist for two more of the constants of motion.
However, determining whether there exists a flux-balance law for $\tilde{Y}$ is far more difficult. In fact, since generically the constraints also evolve (as the constraint depends on the effective metric), determining the average evolution of $\tilde{Y}$ is not even sufficient. The first, and greatest complication, is that the Hamilton propagator no longer has the simple structure that it has in Eq. (121), since solving for the Hamilton propagator requires solving the equations of motion to linear order, which will necessarily introduce constraint terms. Explicitly, Eq. (146b) implies that
@@
$\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau^{\prime}}\left[\delta\tilde{Y}(\tau^{\prime})+V_{\alpha}(\tau,\tau^{\prime})\delta C^{\alpha}(\tau^{\prime})\right]\right\rangle_{\tau^{\prime}}$ $=-[A(\partial_{\upsilon})^{A}+A^{\alpha}(\partial_{\vartheta^{\alpha}})^{A}]\mathcal{F}_{A},$ (154a)
$\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau^{\prime}}\left[U^{\alpha}{}_{\beta}(\tau,\tau^{\prime})\delta C^{\beta}(\tau^{\prime})\right]\right\rangle_{\tau^{\prime}}$ $=[B^{\alpha\beta}(\partial_{C^{\beta}})^{A}-D^{\alpha}(\partial_{\upsilon})^{A}-D^{\alpha\beta}(\partial_{\vartheta^{\beta}})^{A}]\mathcal{F}_{A}+B^{\alpha\beta}\frac{\partial}{\partial C^{\beta}}\langle\mathcal{H}(X,\tau^{\prime})\rangle_{\tau^{\prime}},$ (154b)
@@
This equation shows two important features: first, that the question “what is the average rate of change of $\tilde{Y}$ and $C^{\alpha}$” is not covariant on phase space: one should instead ask questions about the left-hand sides of these equations. It should, however, be noted that it is not clear if these quantities even remain finite, particularly in the case of Eq. (154b) [Eq. (154a) is saved by the fact that the right-hand side is clearly finite]. We will leave to future work the question of whether the quantities on the left-hand side of these equations are useful for practical calculations. The second important feature is that not all terms on the right-hand side are purely asymptotic: they depend on the average value of the coincidence Hamiltonian. If this remains true with a more careful exploration of this system, then this equation is not a “true” flux-balance law.
@@ VI Discussion
Our results here, however, should not be considered as a no-go theorem for true flux-balance laws for spinning bodies in the Kerr spacetime. We present here some options for potentially recovering flux-balance laws, to be explored in future work. First, our results, at this stage, are not entirely explicit: in this paper, the results have mostly depended on the structure of many of the equations, and not in the particular expressions which appear, for example in the various quantities appearing on the right-hand sides of Eqs. (145), (146b), and (146c). It is possible that, once an explicit calculation is completed, many of the offending terms in Eq. (154) will vanish, or simply be higher order in spin. This is suggested by preliminary results showing that $\tilde{Y}$, or a quantity trivially related to it, does not evolve on average in the linear-in-spin approximation [^13] The author thanks Josh Mathews for sharing these preliminary results..
While this could potentially resolve issues with the evolution of $\tilde{Y}$, one remaining issue would be in determining the evolution of the constraint violation terms $C^{\alpha}$. Note that these terms arise due to the fact that we are using the Tulczyjew-Dixon spin supplementary condition associated with the effective metric, but measuring the deviations from this condition as computed from the background metric. Using this choice of spin supplementary condition is the most natural, from the perspective of the generalized equivalence principle. Moreover, it is a necessary ingredient in deriving the Hamiltonian alignment condition and therefore our main result in Eq. (1). However, it may be easier to use the background Tulczyjew-Dixon spin supplementary condition, and simply amend our result accordingly. In this case, the constraint violation terms should not appear. Similarly, it may be useful to consider the Hamiltonian formulations for alternative spin supplementary conditions considered in [46].
@@
Finally, a future direction for this work is to extend the results to second order in $\varepsilon$. Here, there is a major conceptual issue with the calculation in this paper: the first order self force as discussed in this paper is first-order in the sense of a Taylor expansion in powers of $\varepsilon$. However, it is known [42] that such an expansion breaks down on long times: even if the exact trajectory $\Gamma(\varepsilon)$ is initially close to the background worldline $\Gamma$, it will in general diverge on long timescales. This is the motivation for using a self-consistent approach to gravitational self-force, where one uses the exact trajectory $\Gamma(\varepsilon)$ as the source for the metric perturbation, and self-consistently solves for both order-by-order. The key difference between this approach and usual perturbative expansions is that the exact trajectory is never expanded, and so the coefficients at each order in $\varepsilon$ are no longer $\varepsilon$-independent.

### v4

@@ V.2 Spinning particles
which shows that $\Pi^{AB}$ is degenerate; as remarked above in Sec. II.1, this poses no issue. Similarly, there is another Casimir invariant $C_{\star}$ defined by [33]
$C_{\star}\equiv\frac{1}{8}\hat{\epsilon}_{\hat{\alpha}\hat{\beta}\hat{\gamma}\hat{\delta}}\hat{S}^{\hat{\alpha}\hat{\beta}}S^{\hat{\gamma}\hat{\delta}},$ (148)
where $\hat{\epsilon}_{\hat{\alpha}\hat{\beta}\hat{\gamma}\hat{\delta}}$ is the Levi-Civita symbol. Note that $C_{\star}$ vanishes on the physical background phase space as well, and in fact by a suitable choice of coordinates $C_{\star}$ can be chosen to be one of the components of $d^{a}$. As such, explicitly, the components of $C^{\mu}$ are given by $C_{\circ}$, $C_{\star}$, and the two remaining independent components of $d^{a}$.
@@
In the above paragraph, we have made no assumption about the splitting of the five angle variables $\Theta^{\clubsuit}$ into $\vartheta^{a}$ and $\upsilon$. One final assumption which we make is that this angle variable is such that the retarded metric perturbation $h_{ab}$, in terms of its dependence on initial angle variables, can be split into two pieces,
$h_{ab}=\sum_{\boldsymbol{n}}h^{\boldsymbol{n}}_{ab}e^{in_{\alpha}\vartheta^{\alpha}}+\sum_{\boldsymbol{n},N}h^{\boldsymbol{n},N}_{ab}e^{i(n_{\alpha}\vartheta^{\alpha}+N\upsilon)},$ (159)
where $h^{\boldsymbol{n},N}_{ab}$ is $O(S)$. This is the case, for example, in the multiscale expansion in [43]. As such, we have that
$(\partial_{\upsilon})^{A}\mathcal{F}_{A}=O(S^{2}),$ (160)
and so terms proportional to $(\partial_{\upsilon})^{A}\mathcal{F}_{A}$ can be neglected.
We now discuss the consequences of the existence of these coordinates. First, from Eqs. (156) and (158a), it follows that Eq. (2) still holds:
$\left\langle\dot{P}_{\alpha}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=-A^{\beta}{}_{\alpha}(\partial_{\vartheta^{\beta}})^{A}\mathcal{F}_{A}.$ (161)
This implies that one can still write down flux-balance laws for the $P_{\alpha}$, which in this case are given by $m^{2}$, $E$, $L_{z}$, and the generalization of the Carter constant $K$. While such flux-balance laws have been known for $E$ and $L_{z}$ [35, 16], this shows that straightforward generalizations exist for two more of the constants of motion. Note that, despite appearing to have the same form as Eq. (2), the matrix $A^{\alpha}{}_{\beta}$ which appears in Eq. (161) can depend on spin, and so the explicit expressions will not be exactly the same. Splitting the action variables $\mathcal{J}_{\clubsuit}$ into action variables $J_{\alpha}$ and $\Upsilon$, we have that
$A^{\alpha}{}_{\beta}=\frac{\partial P_{\beta}}{\partial J_{\alpha}},$ (162)
as before in Eq. (129). While an explicit full set of action variables has recently been defined in [44], we leave explicit computations of $A^{\alpha}{}_{\beta}$, and therefore an explicit form of the flux-balance law, to future work [43].
That we were able to derive Eq. (161) is due to the “miraculous” coincidence that both $m^{2}$ and $K$, like $E$ and $L_{z}$ but unlike $\tilde{Y}$, are still conserved even when the spin-supplementary condition does not hold. Determining whether there exists a flux-balance law for $\tilde{Y}$ is far more difficult. In fact, since generically the constraints also evolve, determining the average evolution of $\tilde{Y}$ is not even sufficient for determining the total, average motion. To see why the constraints evolve, note that while $\breve{C}^{\alpha}(\varepsilon)$ (that is, the constraints defined using the effective metric) vanishes when evaluated on the exact phase space trajectory $\Gamma(\varepsilon)$, and $C^{\alpha}$ (the constraints in the background metric) vanishes when evaluated on the background trajectory $\Gamma$, the quantity $\delta C^{\alpha}$ is defined by
$\delta C^{\alpha}(\tau)\equiv\left.\delta\Gamma^{A}\nabla_{A}C^{\alpha}\right|_{\Gamma(\tau)}=\left.\frac{\partial C^{\alpha}[\Gamma(\tau;\varepsilon)]}{\partial\varepsilon}\right|_{\varepsilon=0},\vspace{\belowdisplayskip}$ (163)
which does not vanish; it would only vanish if $C^{\alpha}$ on the right-hand side were replaced with $\breve{C}^{\alpha}$.
Returning to the flux-balance law for the Rüdiger constant, the first, and greatest complication, is that the Hamilton propagator no longer has the simple structure that it has in Eq. (132), since solving for the Hamilton propagator requires solving the equations of motion to linear order, which will necessarily introduce constraint terms. Explicitly, Eq. (158b) implies that
@@
$\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau^{\prime}}\left[\delta\tilde{Y}(\tau^{\prime})+V_{\alpha}(\tau,\tau^{\prime})\delta C^{\alpha}(\tau^{\prime})\right]\right\rangle_{\tau^{\prime}}$ $=-A^{\alpha}(\partial_{\vartheta^{\alpha}})^{A}\mathcal{F}_{A},$ (170a)
$\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau^{\prime}}\left[U^{\alpha}{}_{\beta}(\tau,\tau^{\prime})\delta C^{\beta}(\tau^{\prime})\right]\right\rangle_{\tau^{\prime}}$ $=[B^{\alpha\beta}(\partial_{C^{\beta}})^{A}-D^{\alpha\beta}(\partial_{\vartheta^{\beta}})^{A}]\mathcal{F}_{A}+B^{\alpha\beta}\frac{\partial}{\partial C^{\beta}}\langle\mathcal{H}(X,\tau^{\prime})\rangle_{\tau^{\prime}},$ (170b)
@@
This equation shows two important features: first, that the question “what is the average rate of change of $\tilde{Y}$ and $C^{\alpha}$” is not covariant on phase space: one should instead ask questions about the left-hand sides of these equations. It should, however, be noted that it is not clear if these quantities even remain finite, particularly in the case of Eq. (170b) [Eq. (170a) is saved by the fact that the right-hand side is clearly finite]. We will leave to future work the question of whether the quantities on the left-hand side of these equations are useful for practical calculations. The second important feature is that not all terms on the right-hand side are purely asymptotic: they depend on the average value of the conservative Hamiltonian. If this remains true with a more careful exploration of this system, then this equation is not a “true” flux-balance law.
@@ VI Discussion
Our result here, however, should not be considered as a no-go theorem for true flux-balance laws for spinning bodies in the Kerr spacetime. We present here some options for potentially recovering flux-balance laws, to be explored in future work. First, our results, at this stage, are not entirely explicit: in this paper, the results have mostly depended on the structure of many of the equations, and not in the particular expressions which appear, for example in the various quantities appearing on the right-hand sides of Eqs. (157), (158b), and (158c). It is possible that, once an explicit calculation is completed, many of the offending terms in Eq. (170) will vanish, or simply be higher order in spin. This is suggested by recent results showing that $\tilde{Y}$, or rather a quantity trivially related to it (the “parallel component of the spin vector”), does not evolve on average in the linear-in-spin approximation [23, 43]. That is, while the results of this paper imply that there may not be a meaningful sense in which $\tilde{Y}$ possesses a flux-balance law, it may not matter.
While this could potentially resolve issues with the evolution of $\tilde{Y}$, one remaining issue would be in determining the evolution of the constraint violation terms $C^{\alpha}$. Note that these terms arise due to the fact that we are using the Tulczyjew-Dixon spin supplementary condition associated with the effective metric, but measuring the deviations from this condition as computed from the background metric. Using this choice of spin supplementary condition is the most natural, from the perspective of the generalized equivalence principle. Moreover, it is a necessary ingredient in deriving the Hamiltonian alignment condition and therefore our main result in Eq. (1). However, it may be easier to use the background Tulczyjew-Dixon spin supplementary condition, and simply amend our result accordingly. In this case, the constraint violation terms should not appear. Similarly, it may be useful to consider the Hamiltonian formulations for alternative spin supplementary conditions considered in [62]. Since the choice of spin-supplementary condition corresponds only to an arbitrary choice of center-of-mass worldline for an extended body (in particular, there exist formulations which do not require an explicit choice [14]), it is plausible that the asymptotic metric perturbation may be independent of $C^{\alpha}$, and so its averaged evolution is not required.
@@
Finally, a future direction for this work is to extend the results to second order in $\varepsilon$. Here, there is a major conceptual issue with the calculation in this paper: the first order self-force as discussed in this paper is first-order in the sense of a Taylor expansion in powers of $\varepsilon$. However, it is known [53] that such an expansion breaks down on long times: even if the exact trajectory $\Gamma(\varepsilon)$ is initially close to the background worldline $\Gamma$, it will in general diverge on long timescales. This is the motivation for using a self-consistent approach to gravitational self-force, where one uses the exact trajectory $\Gamma(\varepsilon)$ as the source for the metric perturbation, and self-consistently solves for both order-by-order. The key difference between this approach and usual perturbative expansions is that the exact trajectory is never expanded, and so the coefficients at each order in $\varepsilon$ are no longer $\varepsilon$-independent.